Find the length of the rectangle in terms of x.
Area = 4x^3-5x^2-10x-3 and width = 4x+3
Area of rectangle is L times W
This is the last problem for Algebra 1 course, your website helped me with all the other problems, but I can't solve this one. Thank you so much!

Answer by checkley77(12844)

Area = 4x^3-5x^2-10x-3 and width = 4x+3
(4x^3-5x^2-10x-3)/(4x+3)=(x^2-2x-1) is the length.
This requires long division which is very clumbersome to present here.
Proof:
x^2-2x-1
4x+3 (multiply)
-------------------------
4x^3-8x^2-4x+3x^2-6x-3
4x^3-5x^2-10x-3